# RetrievalAttention: Accelerating Long-Context LLM Inference via Vector Retrieval

### v2

@@ Abstract
Transformer-based Large Language Models (LLMs) have become increasingly important. However, due to the quadratic time complexity of attention computation, scaling LLMs to longer contexts incurs extremely slow inference latency and high GPU memory consumption for caching key-value (KV) vectors. This paper proposes RetrievalAttention, a training-free approach to both accelerate attention computation and reduce GPU memory consumption. By leveraging the dynamic sparsity of attention mechanism, RetrievalAttention proposes to use approximate nearest neighbor search (ANNS) indexes for KV vectors in CPU memory and retrieves the most relevant ones with vector search during generation. Unfortunately, we observe that the off-the-shelf ANNS indexes are often ineffective for such retrieval tasks due to the out-of-distribution (OOD) between query vectors and key vectors in attention mechanism. RetrievalAttention addresses the OOD challenge by designing an attention-aware vector search algorithm that can adapt to the distribution of query vectors. Our evaluation shows that RetrievalAttention only needs to access 1–3% of data while maintaining high model accuracy. This leads to significant reduction in the inference cost of long-context LLMs with much lower GPU memory footprint. In particular, RetrievalAttention only needs a single NVIDIA RTX4090 (24GB) for serving 128K tokens in LLMs with 8B parameters, which is capable of generating one token in 0.188 seconds.
@@ 1 Introduction
Recent transformer-based Large Language Models [1] have shown remarkable capabilities in processing long contexts. For instance, Gemini 1.5 Pro [2] has supported the context window of up to 10 million tokens. While this is promising for analyzing extensive data, supporting longer context windows also introduces challenges for inference efficiency due to the quadratic complexity of attention computation. To enhance efficiency, KV caching, a technique that retains past Key and Value vectors, has been widely adopted to prevent redundant computations. However, KV caching-based systems face two primary issues: (a) substantial GPU memory requirements particularly for long contexts, e.g., the Llama-2-7B model requires approximately 500GB per million tokens in FP16 format; and (b) inference latency increases linearly to the context size, primarily due to the time needed to access cached tokens — a common issue across various computing devices, including GPUs. Therefore, reducing token access and storage costs is vital for enhancing inference efficiency.
The solution lies in leveraging the dynamic sparsity inherent in the attention mechanism [3]. This refers to the phenomenon where each query vector significantly interacts with only a limited subset of key and value vectors, with the selection of these tokens varying dynamically based on individual queries. Prior work [4, 5, 6, 7, 8, 9] has proposed various techniques to capitalize on this observation to improve the efficiency of attention computation. However, most of these methods identify important tokens either statically [10, 11] or heuristically [6, 7, 5], leading to imprecise approximations that often result in a significant performance drop. We observe that the Approximate Nearest Neighbor Search (ANNS) index is particularly effective in this context. When using the inner product as the similarity measurement, ANNS precisely aligns with the attention mechanism, represented as $\mathbf{q}\mathbf{K}^{T}$. It can directly identify the most critical key vectors with maximum inner product value, yielding a higher accuracy compared to previous static or heuristic methods (as illustrated in Figure 1).
@@
62.5 125 250 500
Table 1: Decoding latency (on one A100) and GPU memory required for KV cache of Llama-2-7B across different context lengths.
Leveraging ANNS for attention mechanisms presents a unique challenge: the out-of-distribution (OOD) problem between query and key vectors. Most ANNS engines operate under the assumption that both query and key vectors are drawn from the same distribution. However, due to the different projection weights for query and key vectors in attention mechanism, this assumption often does not hold in this context. Unfortunately, the effectiveness of ANNS degrades significantly if the query vectors are OOD. In particular, our empirical analysis shows that maintaining an acceptable accuracy level for inference quality can require scanning 30% of all key vectors to identify the critical ones, which fails to exploit the inherent sparsity in attention mechanism. To the best of our knowledge, this is the first work that identifies the challenge of OOD in using ANNS for attention computation.
In this work, we present RetrievalAttention, an efficient method for accelerating long-context LLM generation. RetrievalAttention employs dynamic sparse attention during token generation, allowing the most critical tokens to emerge from the extensive context data. To address the challenge of OOD, RetrievalAttention proposes a vector index tailored for the attention mechanism, focusing on the distribution of queries rather than key similarities. This approach allows for traversal of only a small subset of key vectors (1% to 3%), effectively identifying the most relevant tokens to achieve accurate attention scores and inference results. In addition, RetrievalAttention reduces GPU memory consumption by retaining a small number of KV vectors in the GPU memory following static patterns (e.g., similar to StreamingLLM [10]) and offloading the majority of KV vectors to CPU memory for index construction. During token generation, RetrievalAttention efficiently retrieves critical tokens using vector indexes on the CPU and merges the partial attention results from both the CPU and GPU. This strategy enables RetrievalAttention to perform attention computation with reduced latency and minimal GPU memory footprint.
We evaluate the accuracy and efficiency of RetrievalAttention on both commodity GPUs (4090) and high-end GPUs (A100) on three long-context LLMs across various long-context benchmarks like $\infty$-Bench [12] and RULER [13]. For the 128K context on the 4090 GPU, RetrievalAttention achieves $4.9\times$ and $1.98\times$ decoding-latency reduction compared to the retrieval method based on exact KNN and traditional ANNS indexing respectively, while maintaining the same accuracy as full attention. To the best of our knowledge, RetrievalAttention is the first solution that supports running 8B-level models on a single 4090 GPU (24GB) with acceptable latency and almost no accuracy degradation.
@@ 2.1 LLM and Attention Operation
In the generation of the $t$-th token, the attention mechanism computes the dot product between the query vector $\mathbf{q}_{t}\in\mathbb{R}^{1\times d}$ (where $d$ is the hidden dimension) and the key vector of past tokens $\mathbf{k}_{i}\in\mathbb{R}^{1\times d}$ (for $i\leq t$). This product is scaled by $d^{-\frac{1}{2}}$ and normalized via a Softmax function to yield the attention score $a_{t,i}$. These scores then weight the values $\mathbf{v}_{i}$, resulting in the output $\mathbf{o}_{t}$.
@@
LLM inference contains two stages: the prefill phase and decoding phase. The prefill phase, which only happens once, computes the keys and values of the prompt with a time-complexity $O(n^{2})$. In the decoding (token generation) phase, the newly generated token becomes the new query and computes attention scores with all past key vectors. One common optimization to avoid repetitive calculation is to cache past KV states in the GPU memory, thereby reducing the complexity to $O(n)$.
@@ 2.2 Expensive Long-Context Serving
Due to the quadratic time complexity of attention operation, serving long-sequence input incurs extremely high cost. Table 1 shows the inference latency without KV cache. When the prompt length reaches 1 million tokens, generating every token requires 1,765 seconds with over 96% of latency spent on attention operations. Although KV cache can reduce the decoding latency, it demands a huge amount of GPU memory for long contexts. As shown in Table 1, 500 GB memory is necessary for storing the KV cache when the context length reaches 1 million tokens, which is far beyond the GPU memory capacity of a single A100 GPU (80 GB). Offloading and reloading the KV cache between GPU and CPU memory is a potential solution but incurs excessive commutation overhead over PCIe [4], degrading the inference performance especially on commodity GPUs.
@@ 2.3 Dynamic and Sparse Attention
Figure 2: (a) Attention is highly sparse in 64,000 tokens. (b) Off-the-shelf ANNS indexes perform poorly on $Q$ to $K$ searches but performs well on $K$ to $K$ searches. The Q and K are dumped from Llama-3-8B with a prompt length of 128,000 tokens. (c) Query vectors ($Q$) are distant from key vectors ($K$) while key vectors ($K$) themselves are close.
Despite the large size of the context, only a small proportion of tokens actually dominate the generation accuracy. Figure 2(a) shows the distribution of $|a_{t,i}|$ in Equation 1 for a query vector from Llama-2-7B with a prompt of 64,000 tokens. We observe that the top 500 tokens dominate the values of $|a_{t,i}|$, while the remaining tokens contribute approximately zero. A high attention score indicates that two vectors are close, as measured by the inner product. Therefore, the sparsity of attention scores means that the relevant keys to the query are very few. We measure the mean-square-error (MSE) of the attention output if we use the top-$k$ $|a_{t,i}|$ as an approximation. We find that it only needs 36 tokens to achieve a very low MSE ($10^{-6}$) of the full attention, showing a high sparsity ratio ($99.9\%$).
Moreover, we observe that as the LLM continues generating new tokens, critical key vectors change dynamically. This means that tokens considered important in previous queries may not be critical in subsequent queries, and vice versa. The dynamic sparsity shows a promising path to approximately compute attention with greatly reduced cost and without sacrificing the model accuracy. For each query, if we can accurately determine the relevant key-value vectors with higher importance, minimum GPU memory and a much lower time complexity can be achieved for attention computation.
@@ 2.4 Challenges of Off-the-shelf Vector Search
Finding the most similar vectors using ANNS indexes is a widely studied problem [14, 15], which semantically aligns with the goal of attention to find the nearest key vectors to each query vector in the inner product space. However, we find that naively applying off-the-shelf vector indexes fails to provide good performance because of the OOD issue between query ($Q$) and key vectors ($K$).
@@
In conventional vector databases, the distribution of vectors between content and query is often well-aligned because they are derived from the same embedding model. However, naively using off-the-shelf vector indexes for attention computation suffers from an inherent distribution gap between queries and keys, which are projected by different model weights. Figure 2(b) (focus on $Q$ to $K$ for now) demonstrates the performance of widely-used vector indexes supported by Faiss [16] using a query vector to retrieve the most similar key vectors. It compares the percentage of keys scanned and the corresponding recall achieved (i.e., the overlapping ratio between the retrieved top-100 results and the ground truth). IVF [14] requires scanning $\sim 30\%$ data for a recall rate higher than 0.95, and HNSW[15] falls into a local optimum. The results show that traditional vector indexes require scanning a large number of vectors to achieve high recall, highlighting the challenge of performing efficient vector searches for attention.
@@
Fundamentally, the difficulty is due to the OOD between query and key vectors. We quantify this using Mahanobis distance [17], which measures the distance from a vector to a distribution. We sample 5,000 vectors from $Q$ and $K$ respectively as the query set and compute the the Mahanobis distance from the query set to the remaining vectors in $K$. Figure 2(c) shows that the queries from $Q$ are significantly distant from the $K$ vectors (OOD) while $K$ themselves are very close. Therefore, traditional index building based solely on the closeness between key vectors does not align with the attention mechanism, which requires to retrieve critical tokens as nearest neighbors from the query vectors’ viewpoint. In contrast, Figure 2(b) shows that using sampled $K$ as the queries ($K$ to $K$) can easily achieve a high recall by only scanning 1-5% vectors because they are in the same distribution. Similarly, query vectors also follow the same distribution. For efficient vector search, the index must consider the OOD characteristic of the attention computation by design.
@@ 3 RetrievalAttention Design
In this work, we focus on the acceleration of token generation and assume the prefill of the long-context prompt is done in advance, which is widely supported by existing LLM service providers (e.g., context caching [18] or separation of prefill and decoding [19, 20]).
@@
We propose RetrievalAttention that leverages attention-aware vector search to accurately approximate attention computation by CPU-GPU co-execution. 3(a) shows the overall design of RetrievalAttention. Based on our observation in §2.3, We derive an approximated attention by selectively retrieving relevant key-value vectors while discarding those that are negligible(§3.1). To efficiently supports long context, we offload most KV vectors to the CPU memory, build vector indexes, and use attention-aware vector search to find critical tokens. (§3.2). To better exploit the GPU devices, we leverage the attention scores obtained in the prefill phase to select a proportion of KV cache that are consistently important during the decoding phase and persist them on GPU devices. RetrievalAttention computes partial attention with dynamically retrieved from CPU memory and persistent key-value vectors in GPU memory in parallel, and finally combine them together(§3.3).

### v3

@@ Abstract
Transformer-based Large Language Models (LLMs) have become increasingly important. However, due to the quadratic time complexity of attention computation, scaling LLMs to longer contexts incurs extremely slow inference speed and high GPU memory consumption for caching key-value (KV) vectors. This paper proposes RetrievalAttention, a training-free approach to both accelerate attention computation and reduce GPU memory consumption. By leveraging the dynamic sparsity of attention mechanism, RetrievalAttention proposes to build approximate nearest neighbor search (ANNS) indexes for KV vectors in CPU memory and retrieve the most relevant ones through vector search during generation. Unfortunately, we observe that the off-the-shelf ANNS indexes are often ineffective for such retrieval tasks due to the out-of-distribution (OOD) between query vectors and key vectors in the attention mechanism. RetrievalAttention addresses the OOD challenge by designing an attention-aware vector search algorithm that can adapt to the distribution of query vectors. Our evaluation demonstrates that RetrievalAttention achieves near full attention accuracy while only requiring access to 1–3% of the data. This leads to a significant reduction in the inference cost of long-context LLMs, with a much lower GPU memory footprint. In particular, RetrievalAttention only needs a single NVIDIA RTX4090 (24GB) to serve 128K tokens for LLMs with 8B parameters, which is capable of generating one token in 0.188 seconds.
@@ 1 Introduction
Recent transformer-based Large Language Models (Vaswani et al., 2017) have shown remarkable capabilities in processing long contexts. For instance, Gemini 1.5 Pro (Team, 2024) has supported the context window of up to 10 million tokens. While this is promising for analyzing extensive data, supporting longer context windows also introduces challenges for inference efficiency due to the quadratic complexity of attention computation. To enhance efficiency, KV caching, a technique that retains past key and value vectors, has been widely adopted to prevent redundant computations. However, KV caching-based systems face two primary issues: (a) substantial GPU memory requirements, particularly for long contexts, e.g., the Llama-3-8B model requires approximately 125GB per million tokens; and (b) inference latency increases linearly to the context size, primarily due to the time needed to access cached tokens — a common issue across various computing devices, including GPUs. Therefore, reducing storage costs and token access is vital for enhancing inference efficiency.
The solution lies in leveraging the dynamic sparsity inherent in the attention mechanism (Deng et al., 2024). This refers to the phenomenon where each query vector significantly interacts with only a limited subset of key and value vectors, with the selection of these critical vectors varying dynamically based on individual queries. Prior work (Tang et al., 2024; Xiao et al., 2024a; Ribar et al., 2024; Lee et al., 2024; Singhania et al., 2024) has proposed various techniques to capitalize on this observation to improve the efficiency of attention computation. However, most of these methods identify important tokens either statically (Xiao et al., 2024b; Li et al., 2024) or heuristically (Xiao et al., 2024a; Ribar et al., 2024; Tang et al., 2024), leading to imprecise approximations that often result in a significant performance drop.
We observe that the Approximate Nearest Neighbor Search (ANNS) index, such as proximity graph (Malkov & Yashunin, 2018), is particularly effective in this context. ANNS index is used to efficiently find the most similar vectors to the query and is widely adopted in various domains like information retrieval (Xiong et al., 2021) and recommendation systems (Cost & Salzberg, 1993; Covington et al., 2016; Pal et al., 2020). When using the inner product as the similarity measurement to build the index for key vectors, searching over the index with the query vector exactly aligns with the attention mechanism.***Maximum inner product search can be viewed as similarity search and efficiently solved by ANNS indexes (Morozov & Babenko, 2018). It can directly identify the most critical key vectors with the maximum inner product to the query vector in sub-linear time complexity, yielding a higher accuracy compared to previous static or heuristic methods (as illustrated in Figure 1). Furthermore, most ANNS algorithms are compatible with CPU implementation, which enables strategic allocation of GPU and CPU memory resources and thus facilitates the handling of longer context inference on devices with limited GPU memory.
@@
15.6 31.2 62.5 125
Table 1: Decoding latency and memory required for KV cache of Llama-3-8B across different context lengths on one A100 GPU.
[…]
In this work, we present RetrievalAttention, an efficient method for accelerating long-context LLM generation. RetrievalAttention employs dynamic sparse attention during token generation, allowing the most critical tokens to emerge from the extensive context data. To address the challenge of OOD, RetrievalAttention proposes a vector index tailored for the attention mechanism, focusing on the distribution of queries rather than keys. This approach allows for the traversal of only a small subset of key vectors (1–3%) to identify the most relevant tokens, yielding accurate attention scores and inference accuracy. In addition, RetrievalAttention reduces GPU memory consumption by retaining a small number of KV vectors in GPU memory following static patterns (e.g., similar to StreamingLLM (Xiao et al., 2024b)) and offloading the majority of KV vectors to CPU memory for index construction. During token generation, RetrievalAttention efficiently retrieves critical tokens using ANNS index on the CPU and merges the partial attention results from both the CPU and GPU. This strategy enables RetrievalAttention to perform attention computation with reduced latency and minimal GPU memory footprint.
We evaluate the accuracy and efficiency of RetrievalAttention on both commodity GPUs (RTX4090) and high-end GPUs (A100) on three long-context LLMs across various long-context benchmarks like $\infty$-Bench (Zhang et al., 2024b) and RULER (Hsieh et al., 2024). For the 128K context on the RTX4090 GPU, RetrievalAttention achieves $4.9\times$ and $1.98\times$ decoding-latency reduction compared to the retrieval method based on exact KNN and traditional ANNS index, respectively, while maintaining the same accuracy as full attention. To the best of our knowledge, RetrievalAttention is the first solution that supports running 8B-level models on a single RTX4090 GPU (24GB) with acceptable latency and almost no accuracy degradation.
@@ 2.1 LLM and Attention Operation
In the generation process of the $t$-th token, the attention operation computes the dot product between the query vector $\mathbf{q}_{t}\in\mathbb{R}^{1\times d}$ (where $d$ is the hidden dimension) and the key vectors of all preceding tokens $\mathbf{k}_{i}\in\mathbb{R}^{1\times d}$ (for $i\leq t$). This product is scaled by $d^{-\frac{1}{2}}$ and normalized via a Softmax function to yield the attention score $a_{t,i}$. These scores then weight the values $\mathbf{v}_{i}$, resulting in the output $\mathbf{o}_{t}$.
@@
LLM inference contains two stages: the prefill phase and decoding phase. The prefill phase, which only happens once, takes all tokens of the prompt as input and performs attention with a time-complexity $O(n^{2})$. In the decoding (token generation) phase, the newly generated token is added to the input and computes attention scores with same complexity. One common optimization to avoid repetitive calculation is caching past KV states, thereby reducing the complexity to $O(n)$.
@@ 2.2 Expensive Long-Context Serving
Due to the quadratic time complexity of attention operation, serving long-sequence input incurs extremely high costs. Table 1 shows the inference latency of Llama-3-8B without KV cache. When the prompt length reaches 1 million tokens, generating every token requires 1,765 seconds with over 96% of latency spent on attention operations. Although KV cache can reduce the decoding latency, it demands a huge amount of GPU memory for long contexts. As shown in Table 1, 125 GB memory is necessary for storing the KV cache when the context length reaches 1 million tokens, which is far beyond the GPU memory capacity of commodity GPUs such as the RTX4090 (24GB) or even high-end GPUs like A100 (40GB or 80GB). This necessitates either scaling to more GPUs to accommodate the large KV cache  (Liu et al., 2024a) or repetitively offloading and reloading the entire KV cache between CPU and GPU memory over PCIe  (Sheng et al., 2023), resulting in excessive communication overhead. Neither approach provides an efficient and cost-effective solution for long-context inference on commodity GPUs.
@@ 2.3 Dynamic and Sparse Attention
Figure 2: The dynamic sparsity of each layer and head in Llama-3-8B model in the KV retrieval test of 100,000 tokens. The blue curve shows that using dynamically selected top-1000 critical tokens achieves an average recovery ratio of 89%, indicating high attention sparsity. In contrast, the orange curve reveals that statically using the initially determined top-1000 critical tokens from the generation of the first token to generate subsequent tokens drops the average recovery ratio to 71%.
Corroborating recent work (Xiao et al., 2024b; Li et al., 2024), we observe that attention computation in LLMs exhibits significant sparsity. Despite the large context length, only a small fraction of tokens with the highest attention scores (i.e., $a_{t,i}$ in Equation 1), also known as critical tokens, contribute significantly to the attention output.
We quantify the attention sparsity by calculating the cumulative sum of attention scores of top-$k$ critical tokens. This cumulative sum, called recovery ratio, represents how much of the full attention can be recovered using a small number of critical tokens, with a higher recovery ratio indicating greater sparsity. When generating 20 tokens consecutively based on a prompt of 100,000 tokens, we profile the average recovery ratio of decoding tokens using top-1000 critical tokens in different layers and heads of the model. As shown in the blue curve of Figure 2, by accurately selecting top-1000 critical tokens based on full attention, most attention heads can recover over 90% of the attention scores from the full attention, with an average of 89% across all heads and layers.
Furthermore, we observe that as the LLM continues generating new tokens, the critical key vectors change dynamically, highly depending on the current query vector. To verify this, we first collect the top-1000 critical key vectors to generate the first token in each attention head and statically apply them for the subsequent token generation. The results shown in the orange curve of Figure 2 indicate a significant drop in the average recovery rate, from 89% to 71%. This demonstrates that tokens considered important in previous queries may not be critical in subsequent queries, and vice versa. Therefore, it is necessary to dynamically select important tokens for each query vector.
The dynamic sparsity shows a promising path to approximately compute attention with greatly reduced cost and without sacrificing the model accuracy. For each query, if we can accurately identify the relevant key-value vectors with higher importance, minimum GPU memory and a much lower time complexity can be achieved for attention computation.
@@ 2.4 Challenges of Off-the-shelf Vector Search
Figure 3: (a) Query vectors ($Q$) and key vectors ($K$) are dumped from Yi-9B and Llama-3-8B with a prompt length of 128,000 tokens. Off-the-shelf ANNS indexes perform poorly on $Q$ to $K$ searches while work well for $K$ to $K$ searches. (b) Query vectors are distant from key vectors, while key vectors themselves are close.
To reduce the latency of long contexts inference while maintaining performance, we require a method to accurately identify the critical tokens to the current query in sub-linear time complexity. Additionally, given the constrained GPU memory, it would be beneficial if such a method could efficiently utilize CPU memory to manage the KV vectors. Based on Equation 1, one key vector is critical for a query vector if they have a large inner product. With the inner product as a similarity function, performing searches on ANNS indexes aligns well with the goal of the attention mechanism to efficiently find critical key vectors for a query vector.
Traditional ANNS indexes generally cluster similar (close) vectors and select the representative vector for each cluster (Sivic & Zisserman, 2003) or directly build connections between similar vectors to form a proximity graph (Wang et al., ).In this context, we use “similar” and “close” to indicate vectors with larger inner product interchangeably. For cluster-based indexes, the query first compares with all representative vectors and then only accesses the most similar clusters, whereas, in the proximity graph, the query performs a greedy search, moving closer to the most similar vectors at each hop. Both methods typically require scanning a limited subset of all vectors (e.g., 1%) to identify the most similar vectors to the query, achieving high search efficiency and accuracy. However, we find that naively applying off-the-shelf vector indexes fails to provide good performance because of the OOD issue between query ($Q$) and key vectors ($K$).
@@
In conventional vector databases, the distribution of vectors between content and query is often well-aligned because they are derived from the same embedding model. However, naively using traditional vector indexes for attention computation suffers from an inherent distribution gap between queries and keys, which are projected by different weights as 2.1. Figure 3(a) (focus on $Q$ to $K$ for now) demonstrates the performance of widely-used vector indexes supported by Faiss (Douze et al., 2024) using a query vector to retrieve the most similar key vectors. It compares the percentage of keys scanned and the corresponding recall achieved (i.e., the overlapping ratio between the retrieved top-100 results and the ground truth). Cluster-based IVF (Sivic & Zisserman, 2003) requires scanning $\sim$30–50% data for a recall rate higher than 0.95, and graph-based HNSW (Malkov & Yashunin, 2018) falls into a local optimum. The results show that traditional vector indexes require scanning a large number of vectors to achieve a high recall, highlighting the challenge of performing efficient vector searches for attention.
@@
Fundamentally, the difficulty is due to the OOD between query and key vectors. We quantify this using Mahanobis distance (Mahalanobis, 2018), which measures the distance from a vector to a distribution. We sample 5,000 vectors from $Q$ and $K$ respectively as the query set and compute the the Mahanobis distance from the query set to the remaining vectors in $K$. Figure 3(b) shows that the queries from $Q$ are significantly distant from the $K$ vectors (OOD) while $K$ themselves are very close. Therefore, traditional index building based solely on the closeness between key vectors does not align with the attention mechanism, which requires to retrieve critical tokens as nearest neighbors from the query vectors’ viewpoint. In contrast, Figure 3(a) shows that using sampled $K$ as the queries ($K$ to $K$) can easily achieve a high recall by only scanning 1–5% vectors because they are in the same distribution. Similarly, query vectors in each attention head also follow the same distribution as they are projected by the same model weight. For efficient vector search, the index must consider the OOD characteristic of the attention computation by design.
@@ 3 RetrievalAttention Design
In this work, we focus on the acceleration of token generation and assume the prefill of the long-context prompt is done in advance, which is widely supported by existing LLM service providers (e.g., context caching (Google Cloud, 2024) or separation of prefill and decoding (Patel et al., 2024; Qin et al., 2024)).
@@
We propose RetrievalAttention that leverages attention-aware vector search to approximate attention computation by CPU-GPU co-execution accurately. 4(a) shows the overall design of RetrievalAttention. Based on our observation in §2.3, We derive an approximated attention by selectively retrieving relevant key-value vectors while discarding those that are negligible(§3.1). To efficiently support long context, we offload most KV vectors to the CPU memory, build vector indexes, and use attention-aware vector search to find critical tokens. (§3.2). To better exploit the GPU devices, we leverage the attention scores obtained in the prefill phase to select a proportion of KV cache that is consistently important during the decoding phase and persist them on GPU devices. RetrievalAttention computes partial attention with dynamically retrieved from CPU memory and persistent key-value vectors in GPU memory in parallel and finally combines them together(§3.3).